OpenStudy (anonymous):

What is the equation in slope intercept form of the line that has a slope of-3 and passes through the point (0,5)

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

You use the point-slope of the line to do this: \[y-y _{1} = m ( x-x _{1} )\] You know m, the slope. \[(x _{1},y _{1})\] represents the co-ordinates of the point on the line that you are also given, in this case (0,5). Just put the values into the equation and you've finished your work.

OpenStudy (anonymous):

Y-8=6(x-1)?

OpenStudy (anonymous):

No. The point is (0,5). y - 5 = -3 ( x - 0 ) Do you see why?

OpenStudy (tkhunny):

Why are we using the Point-Slope Form to solve this? What is the equation in slope intercept form of the line that has a slope of-3 and passes through the point (0,5) We are GIVEN the Slope and the Intercept. y = -3x + 5 Done!

OpenStudy (anonymous):

Ho, ho! Because I didn't notice that it was the intercept.